OpenStudy (coexist2):

A triangle has sides the lengths 18, 80, and 82... Is it a right triangle, explain? A. no, 18 squared+80 squared= 82 squared. (dash through equal sign.) B. yes, 18 squared+80 squared = 82 squared. C. no, 18 squared+80 squared= 82 squared. D. yes 18 squared+ 80 squared= 82 squared. (dash through equal sign.)

OpenStudy (jtvatsim):

You can test whether a triangle is a right triangle by seeing if it meets the Pythagorean Theorem: \[a^2 + b^2 = c^2\] If it is equal, then it must be a right triangle. If it is NOT equal, then it is NOT a right triangle.

OpenStudy (jtvatsim):

For example, is a triangle with sides 3, 4, and 5 a right triangle? Test: \[3^2 + 4^2 = 9 + 16 = 25\] and \[5^2 = 25\] so, \[3^2 + 4^2 = 5^2\] Since they are equal, this is a right triangle! Try it with your numbers now. : )

OpenStudy (coexist2):

thanks for the help everything equals 6724

OpenStudy (wolf1728):

A and D also say "dash through equal sign menaing NOT equal

OpenStudy (jtvatsim):

Good job! So, you can say "YES!" They are equal, so this must be a right triangle.

OpenStudy (coexist2):

its B.

OpenStudy (jtvatsim):

Yes